OpenStudy (caozeyuan):

For a cubic polynomial a3x3+a2x2+a1x+a0, the roots r1, r2, and r3 come in three symmetric combinations: σ1=r1+r2+r3, σ2=r1r2+r1r3+r2r3, and σ3=r1r2r3. The sum of the kth powers of the roots is defined as Sk=r^k1+r^k2+r^k3. a) Prove that a3*S1+a2=0. b) Prove that a3*S2+a2*s1+2*a1=0. c) Generalize to something that starts with a3^Sk. Generalize further to polynomials of degree n.

OpenStudy (caozeyuan):

@experimentX @whpalmer4 help me please!

OpenStudy (caozeyuan):

@mukushla look at this please

OpenStudy (experimentx):

\[ a_3x^3+a_2x^2+a_1x+a_0 = 0 \implies x = r_1, r_2, r_3\]

OpenStudy (experimentx):

what are k's?

OpenStudy (caozeyuan):

any real number can be k

OpenStudy (caozeyuan):

usually integers to make life easier, though

OpenStudy (experimentx):

no ,, there are k1,k2,k3 <-- they are not defined

OpenStudy (caozeyuan):

I think it's related with Newton's Sum in some way

OpenStudy (caozeyuan):

it's actually r1^k+r2^k+r3^k. I made a typo earlier

OpenStudy (experimentx):

nvm got that.

OpenStudy (experimentx):

Do you know Vieta's fomula?

OpenStudy (experimentx):

\[ a_3x^3+a_2x^2+a_1x+a_0 = a_3 (x - r_1)(x-r_2)(x-r_3) \] compare the coefficients of x ... it should work out smoothly.

OpenStudy (caozeyuan):

great, thx! but how about generalize it to polynomials with degree n?

OpenStudy (experimentx):

as you stated this is called newton's sum generalization of proof is here http://digitalcommons.unl.edu/cgi/viewcontent.cgi?article=1049&context=mathfacpub

OpenStudy (experimentx):

interesting proof ... cleverly manipulated

OpenStudy (caozeyuan):

too complicated to understand.......

OpenStudy (experimentx):

which part you don't understand?

OpenStudy (caozeyuan):

All of them! I'm just a high school stu, not a univ mathematician!

OpenStudy (experimentx):

do you know basic stuff like differentiation? fundamental theorem of calc?

OpenStudy (caozeyuan):

Yeap, I've done AP cal BC,so I know all the stuff about it

OpenStudy (experimentx):

Let \[ x^n + a_n x^{n-1} + .. +a_1\] be your polynomial. and \( r_1, r_2, ... r_n\) be it's zeros. from fundamental theorem of calc you have \[ x^n + a_n x^{n-1} + .. +a_1 = (x-r_1) (x-r_2) ... (x-r_n)\]

OpenStudy (caozeyuan):

is that fund. theor. of cal or of alg?

OpenStudy (experimentx):

typo ..algebra sorry.

OpenStudy (experimentx):

differentiate it one ... on LHS, one term is missing. \[ Q(x) = (x-r_1) (x-r_2) ... (x-r_n)\]

OpenStudy (experimentx):

Q'(x) = (x-r2)...(x-rn)+(x-r1)(x-r3) ... (x-rn)+ ....+(x-r1)...(x-r_n-1) we are using product rule.

OpenStudy (experimentx):

divide Q'(x)/Q(x) and see what you get

OpenStudy (caozeyuan):

my head just exploded!